OpenStudy (anonymous):

Rewrite the following functions in intercept form. Find the x−intercepts and the vertex. 1. y = x^2−2x−8 2. y =−x^2+10x−21 3. y = 2x^2+6x+4 4. y = 3(x+5)(x−2)

Parth (parthkohli):

hi, nin

Parth (parthkohli):

that said, do you have an idea of where to start with this?

OpenStudy (anonymous):

this isnt nin.......

OpenStudy (anonymous):

im going to kill him now

Parth (parthkohli):

yeah, okay. do you know what "x-intercept" means?

OpenStudy (anonymous):

I can help with this one too if you want Parth

OpenStudy (anonymous):

.......thanks his gf is stupid....

OpenStudy (zzr0ck3r):

I think you mean vertex form?

Parth (parthkohli):

yes, back to the question now.

OpenStudy (anonymous):

I'm going to assume you know what an x-intercept is, that is where you set your equation to 0

Parth (parthkohli):

no, I think the vertex means a point (x,y) where the function is minimum/maximum.

OpenStudy (anonymous):

vertex is the (x,y) and is found by using -b/2a to get x coordand plug that back into the equation to find the y coord

OpenStudy (zzr0ck3r):

for \(f(x) = x^2+bx+c\) the vertex form is \(f(x) = (x+\frac{b}{2})^2+c-(\frac{b}{2})^2\)

Parth (parthkohli):

in the case of quadratic functions, the vertex is given by\[\left(-b/2a, f(-b/2a)\right)\]or, if you notice\[(-b/2a, -D/4a)\]

Parth (parthkohli):

\[D = b^2 - 4ac\]

OpenStudy (zzr0ck3r):

ok but I dont know what intercept form is, im thinking she was saposed to get it in vertex form, then one can see the vertex....

Parth (parthkohli):

x-intercepts = roots of the function

OpenStudy (anonymous):

yeah, I'm thinking that should be vertex form which is (x-h)^2+k

OpenStudy (zzr0ck3r):

right, what is x intercept form?

Parth (parthkohli):

well, I mean the points at which the roots lie.

OpenStudy (zzr0ck3r):

lol

OpenStudy (zzr0ck3r):

the first line of the question Rewrite the following functions in intercept form

OpenStudy (anonymous):

@msumner, does it ask for intercept form or vertex form?

OpenStudy (zzr0ck3r):

maybe intercept form is factored?

OpenStudy (anonymous):

it says both

OpenStudy (zzr0ck3r):

nah because the last one is factored

OpenStudy (zzr0ck3r):

@msumner reread the question, does it say put it in vertex form and find the vertex and intercepts or does it say put it in intercept form and find the vertex and intercepts?

Parth (parthkohli):

I'm going to demonstrate one such question to you.\[y = x^2 + 10x + 25\] To get the roots, solve \(x^2 + 10x + 25 = 0\). You will get \(x = -5\) as the root. The point(s) at which the roots fall are the x-intercepts. Here, the x-intercept is \((-5,0)\). The vertex is \((-b/2a, f(-b/2a)) \implies (-5,0)\).

OpenStudy (anonymous):

To find the x-intercepts, I would first try to factor out the quadratic and if it doesn't factor, use the quadratic formula

OpenStudy (anonymous):

This one does factor

Parth (parthkohli):

\[y = x^2 + 7x + 12\]Roots: 3, 4 Points/x-intercepts: \((3,0)\) and \((4,0)\). Vertex: \((-7/2, -1)\)

OpenStudy (zzr0ck3r):

I will demonstrate a similar equation in vertex form, (just in case) \(y = x^2+10x+12 = (x+5)^2+12-25 = (x+5)^2-13\)

OpenStudy (anonymous):

Really good examples here, @msumner take notes :)

Parth (parthkohli):

The vertex form in general is\[ax^2 + bx + c = a\left(x + \dfrac{-b}{2a}\right)^2 + \dfrac{4ac - b^2}{4a}\]

OpenStudy (anonymous):

That is standard form I thought

Parth (parthkohli):

Yes, that's the standard form for the vertex form.

OpenStudy (anonymous):

vertex form that I know is (x-h)^2 + k

OpenStudy (nincompoop):

I didn't know parabolas have intercept form :D

OpenStudy (zzr0ck3r):

they are the same thing,... parth just showed how to get there from ax^2+bx+c

OpenStudy (zzr0ck3r):

there is no such form as far as I know....

OpenStudy (zzr0ck3r):

on any function

Parth (parthkohli):

I mean there is\[\left(\dfrac{-b\pm \sqrt{b^2 - 4ac}}{2a}\right)\]

Parth (parthkohli):

and the y-coordinate is zero.

OpenStudy (anonymous):

The intercept form of a quadratic equation is as follows: y=a(x−p)(x−q)

OpenStudy (nincompoop):

there is the standard form and the general (vertex) form that I know of

OpenStudy (anonymous):

I just found that when I was looking it up

Parth (parthkohli):

so... the factored form.

OpenStudy (zzr0ck3r):

I almost guarantee the OP needed to put them in vertex form and then use that form to see the vertex clearly...

OpenStudy (anonymous):

that is what it is called

OpenStudy (zzr0ck3r):

vertex form, completing the square, the quadratic formula are all the same thing...

Parth (parthkohli):

Well, anyway, there's the general form. You may not even feel the need to complete the square. They're the same processes though, so feel free.

OpenStudy (anonymous):

I think it is just faster in general to use the quadratic formula unless I can see right away that it can be factored

OpenStudy (nincompoop):

what is this? y=a(x−p)(x−q) what is the q and what is the p? and what is the x in there?

OpenStudy (nincompoop):

we are throwing bunch of letters LAUGHING OUT LOUD

Parth (parthkohli):

Of course, but the quadratic formula ain't no help with your vertex.

OpenStudy (anonymous):

p and q are the x-intercepts (also zeros, solutions, roots) a is the vertical stretching factor

OpenStudy (anonymous):

there is the context to the intercept form (factored form)

OpenStudy (zzr0ck3r):

if you get fast at completing the square you can solve them faster than you can write down the quadratic formula

OpenStudy (zzr0ck3r):

jt the last one is in factored form

OpenStudy (nincompoop):

always convert from standard to vertex by completing the square if the equation isn't already a completed square

OpenStudy (zzr0ck3r):

so I ruled that out

OpenStudy (skullpatrol):

Any questions @msumner ?

OpenStudy (zzr0ck3r):

all in all I feel bad for the person that asked this question

OpenStudy (zzr0ck3r):

lol @skullpatrol

OpenStudy (nincompoop):

LAUGHING OUT LOUD

OpenStudy (anonymous):

yeah, she got a lot of information just now I hope we don't overwhelm her

Parth (parthkohli):

She started with four questions. She must have double of them now.

OpenStudy (nincompoop):

I am glad I am only in pre-algebra

OpenStudy (zzr0ck3r):

she started writing and just said f**k it

OpenStudy (anonymous):

I am in pre-calculus

OpenStudy (zzr0ck3r):

im in algebra

OpenStudy (zzr0ck3r):

544:)

Parth (parthkohli):

I'm 8 and what is this

OpenStudy (anonymous):

im good lol

OpenStudy (anonymous):

wait, if you are in algebra and pre-algebra, how do you know algebra 2 concepts?

OpenStudy (zzr0ck3r):

im in abstract algebra as a graduate student:)

OpenStudy (anonymous):

thanks guys hahaha..... nin is just confusing you .....

OpenStudy (anonymous):

ah, that explains it :) anyway, hope we helped you out @msumner and sorry for the long explanations

Parth (parthkohli):

nin, why are you referring to yourself in third person?

OpenStudy (zzr0ck3r):

lol

OpenStudy (anonymous):

its fine i got a laugh out of it :)..... im not nin smarty

OpenStudy (anonymous):

http://strawpoll.me/1831262/r

OpenStudy (zzr0ck3r):

adding a y to the end of words makes you sound like a girl

OpenStudy (zzr0ck3r):

its girly

OpenStudy (anonymous):

that is awesome @iambatman, I am adding that to my bookmarks

OpenStudy (anonymous):

Spelt the name wrong, no biggie, just being a nincompoop today.

OpenStudy (anonymous):

im a girl named mickie

OpenStudy (anonymous):

you don't have to tell us your real name

Parth (parthkohli):

*sigh* nin, is that the best name you could come up with?

OpenStudy (anonymous):

who said that was my real name

OpenStudy (zzr0ck3r):

thats a lie, no parent would name a girl mikey

OpenStudy (anonymous):

Unless they're fan of rats

OpenStudy (anonymous):

lol its a nickname

Parth (parthkohli):

of course it's not your real name your real name is what nin's real name is you're nin

OpenStudy (zzr0ck3r):

was that a goonies joke?

OpenStudy (anonymous):

I will let you guys figure out this mystery and I will get back to my hw

OpenStudy (anonymous):

lol